OpenStudy (anonymous):

If antique brass is 30.0% zinc and 70.0% copper, how many grams of antique brass can be made from 18.0 grams of zinc and 35.1 grams of copper?

OpenStudy (anonymous):

can you explain it?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@aditya96

OpenStudy (anonymous):

what do you mean if it's a chemistry question, what would be the difference

OpenStudy (tkhunny):

Your solution will be one of these: 1) Use all the zinc, but not all the copper. 2) Use all the copper, but not all the zink. 3) Use both completely. (This will be just luck.) Let B = The number of ounces of brass that can be created Use all the zinc B*0.30 = 18.0 ==> B = 60 but this requires 60*0.70 = 42 oz of copper, and we've only 35.1 Use all the copper B*0.70 = 35.1 ==> B = 50.14 and this requires 15.04 oz of zinc. I think we're done.

OpenStudy (primeralph):

isn't that what I did?

OpenStudy (tkhunny):

No, @primeralph, you simple dropped some algebra and provided no explanation. Let's try harder to teach.

OpenStudy (anonymous):

you meant to say grams, not ounces, right?

OpenStudy (tkhunny):

Obvious to whom? Obvious or not, we're teaching methods and thought processes, here. It's about helping students to see, not just telling that they should see.

OpenStudy (tkhunny):

Right. Sorry about that. Normally I would check the units quite carefully, but everything was the same in this problem, so I guess I fell asleep on that point. Apologies.

OpenStudy (anonymous):

So my final answer would be 50.14?

OpenStudy (tkhunny):

ahahaha - No! Where are your units?! Didn't you just call me on that? :-)

OpenStudy (anonymous):

Oh man I'm so confused. I'm sorry!

OpenStudy (anonymous):

nevermind I got it!

OpenStudy (anonymous):

No, I'm not going to interrupt, you can continue. But it would be so much simpler if you noticed the ratio

OpenStudy (anonymous):